# CudaForge: An Agent Framework with Hardware Feedback for CUDA Kernel Optimization

### v1

@@ Abstract
5555606065657070757580808585909095951001000.60.60.80.8111.21.21.41.41.61.61.81.8OpenAI-$o3$-self-refineo3-correctiono3-optimizationCudaForgeCudaForgeKevin-32BAgentic BaselineCorrectness (%)Performance (× Speedup)w/o Judgew/o correction feedbackw/o optimization feedbackEvaluated on Level 1–2Evaluated on Level 1–3
Figure 1: CudaForge achieves state-of-the-art results on KernelBench in both correctness and performance, surpassing RL-based methods such as Kevin-32B [1], the agentic baseline [2], and OpenAI-o3 [3]. To further evaluate the effectiveness of our design, we additionally develop three customized variants of OpenAI-o3: o3-self-refine, o3-correction, and o3-optimization, which serve as baselines for ablation comparison. Experimental details are provided in Section 3.
@@ 1 Introduction
Motivation. CUDA has become the de facto standard for deep learning training because modern frameworks such as PyTorch and TensorFlow are deeply integrated with NVIDIA’s optimized GPU libraries [4]. Efficient CUDA kernels are crucial for accelerating deep learning workloads [5, 6] .
However, developing high-efficiency CUDA kernels is known to be challenging with a steep learning curve, requiring deep expertise in GPU architectures and parallel programming [7]. For example, it took more than 2 years from the debut of the Hopper GPU architecture to the release of FlashAttentionV3 [8], which is specially designed for Hopper GPUs.
This high development barrier has driven growing interest in finding automated ways of generating highly efficient and customized CUDA kernels. For example, some work [9] [10] employs auto-tuning and evolutionary search to automatically explore kernel implementation spaces and optimize low-level parameters for specific hardware. More recently, there has been a growing interest in leveraging large language models (LLMs) to perform such tasks. LLM is believed to hold great promise in generating efficient and high-quality kernels, due to its capability of code generation in other domains, such as Python and C++ [11, 12].
Existing Works and Key Challenges. Generally, using LLMs for CUDA kernel generation is still in an early stage. In KernelBench [13], the authors attempt to directly use state-of-the-art (SOTA) models, such as OpenAI-o1 and Claude-3.5-Sonnet, to generate kernels. However, it has been observed that these SOTA models still struggle to produce correct or performant kernels out of the box, revealing fundamental limitations of existing LLMs in this domain.
To address this gap, recent studies have explored two main paradigms. The first approach is based on reinforcement learning (RL) [14, 15]. CUDA-L1 [16] and Kevin [1] adopt RL to enhance LLMs’ ability to generate correct and performant CUDA code.
The second approach is based on AI agents. In particular, in an independent and contemporaneous work [2]111published on arxiv Sept 16th, 2025, researchers have explored agentic frameworks at inference time. Agents project PyTorch methods into CUDA kernel designs, then the CUDA kernels are further refined by sampling new kernels and verification filtering. This design effectively improves correctness in CUDA kernel generation without the high cost of RL training.
@@
(C2) High training and inference cost. RL-based approaches such as [16, 1] require substantial computational resources and long training cycles, making them unsuitable for low-resource or rapid-prototyping settings. In addition, multi-stage agentic pipeline developed by [2] incurs high inference costs (about 6 H100 hours and \$5 API cost per kernel), which greatly limits its practical applicability of the approach.
(C3) Lack of hardware feedback. Human experts typically follow an iterative workflow to develop performant CUDA kernels through testing and refinement. They rely on hardware feedback like Nsight Compute (NCU) [^2] Nsight Compute (NCU) is NVIDIA’s official kernel-level profiler for CUDA programs. to identify bottlenecks and optimize kernels accordingly [17, 18, 19]. In contrast, RL-based approaches [16, 1] train LLMs to directly generate or optimize kernels, but do not incorporate hardware feedback at all. As a result, they rely on blind exploration during generation, lacking targeted guidance. This often leads to suboptimal kernel efficiency, limiting their practical applicability.
@@
Our Contributions. To address these challenges, we propose CudaForge, a simple, effective, and low-cost multi-agent workflow for CUDA kernel generation and optimization, as shown in Figure 2. Our workflow is inspired by the iterative workflow of human experts [17, 18, 19], which contains steps such as developing initial kernels, testing correctness, analyzing hardware feedback, and iterative improvement.
@@
Compared to single-LLM approaches that generate and evaluate code using the same LLM, our framework separates these roles into an independent Coder and Judge, enabling more specialized reasoning and more reliable iterative refinement. Unlike RL-based methods, CudaForge is training-free, avoiding the substantial cost of policy training. It is also hardware-aware, allowing it to tailor CUDA kernel optimizations to the underlying system, making the proposed framework easily generalizable across different GPUs. Finally, in contrast to existing multi-agent frameworks [2], CudaForge is lightweight and cost-efficient, running in just 26.5 minutes on a single RTX 6000 GPU and about \$ 0.3 per kernel in API on average, while still achieving significantly better performance.
We evaluate CudaForge on 250 KernelBench tasks from Level 1 to Level 3. Though these tasks are challenging, CudaForge attains a 97.6% correctness rate and delivers an average speedup of 1.68× over PyTorch baselines, which significantly outperforms advanced RL model like Kevin-32B, advanced frontier model like OpenAI-o3 [3] and Agentic Baseline [2], shown in Figure 1. Further, we have conducted comprehensive ablation studies of the features of CudaForge, such as its effectiveness across multiple GPU architectures, its inference-time scalability by increasing the number of generation, and the effect of different base models. Overall, we observed that the proposed CudaForge achieves robust performance in all these settings.
@@ 2.1 CudaForge Framework
Inspired by the iterative workflow of human experts [17, 18, 19], we design CudaForge as an iterative multi-agent framework, illustrated in Figure 2. The framework involves two independent agents: a Coder and a Judge. The Coder generates candidate kernels based on the task description and feedback from the Judge, while the Judge evaluates each candidate using the kernel itself, hardware feedback, and runtime information.
@@ 2.2 Component Design
The Coder acts as the generative module in our framework, responsible for producing functionally correct and efficient CUDA kernel candidates given the task description, previous candidates, and feedback from the Judge. Directly generating a CUDA kernel from scratch is a highly challenging task for LLMs [13]. Following the setting of KernelBench [13], we provide the Coder with a one-shot demonstration in the prompt, which includes an example pair of a PyTorch reference and its corresponding CUDA kernel implementation. This example serves as a structural and stylistic guide, helping the Coder understand the expected syntax, memory access patterns, and API usage required for valid kernel generation.
@@
After the Coder generates a candidate kernel, our workflow proceeds to verify its correctness through a two-stage testing process, including compilation and execution. The compilation stage ensures that the generated kernel is syntactically valid and can be successfully compiled into executable CUDA code. The execution stage then evaluates the functional correctness of the compiled kernel by running it on predefined test cases. We compare the kernel’s outputs with those from the PyTorch reference implementation under the same inputs, and consider the kernel functionally correct only if the numerical difference between the two outputs is within a small tolerance (set to $1\text{e}{-4}$ in our experiments). A kernel is regarded as correct only when it passes both compilation and execution stages successfully, which is a commonly adopted criterion [1, 2, 13].
@@ 3.1 Benchmark and Evaluation
We evaluate our method and baselines on KernelBench [13], a popular benchmark designed to assess the ability of LLMs to generate CUDA kernels. KernelBench consists of multiple difficulty levels, and we adopt all tasks from Level 1 to Level 3, resulting in a total of 250 tasks. Specifically, Level 1 contains relatively simple 100 tasks involving basic operators (e.g., matrix multiplication), Level 2 includes medium-difficulty 100 tasks composed of multi-step operator combinations, and Level 3 contains 50 challenging tasks involving full neural network architectures (e.g., AlexNet). Each task is accompanied by a reference PyTorch implementation and predefined input/output specifications, which enables fully automated and reliable evaluation of both correctness and performance. Details of KernelBench are provided in Appendix D.1.
@@ 3.2 Settings & Baselines
  • Kevin-32B: a strong RL-based model for CUDA kernel optimization from [1]. We directly take results from their official paper.
  • Agentic Baseline: the agentic workflow from [2], a strong multi-agent baseline, which uses an LLM ensemble including both reasoning (o3 & o4-mini) and conventional LLMs (Claude Sonnet 3.7 & GPT-4.1).
@@
Table 1: Main results on KernelBench (Level 1-3, 250 tasks). Results of Agentic Baseline is on Level 1 and 2. All experiments here are run in RTX 6000. Methods evaluated on $\mathcal{D}^{*}$ are marked with .
@@ 3.4 Comparison with Kevin-32B on H200
In this subsection, we compare CudaForge with Kevin-32B [1]. However, since Kevin-32B is not open-sourced—meaning that key details such as test-time prompts, evaluation setup, and benchmark specifications are not publicly available—a fully fair and reproducible comparison cannot be conducted. According to the original paper, Kevin-32B was trained and evaluated on H200 GPUs using a self-constructed benchmark of comparable difficulty to KernelBench Levels 1 and 2. To ensure the comparison is as fair as possible under these constraints, we re-evaluate CudaForge on the same H200 hardware across KernelBench Levels 1–3, aligning as much as possible our computational environment and evaluation protocol with those reported for Kevin-32B.
@@ 3.5 API and Computation Time Cost of CudaForge
Table 3: Comparison of API and computation time cost between CudaForge and the Agentic Baseline [2].
@@
As shown in Table 3, CudaForge requires on average only 26.5 minutes of wall-clock time on a single RTX 6000 GPU and incurs merely \$0.3 of API cost per kernel. This is highly cost-efficient compared with the Agentic Baseline [2], which reports approximately 6 GPU hours on H100 and \$5 per kernel in their Appendix E.
@@
We further analyze how the API and computation costs relate to CudaForge’s performance, as illustrated in Figure 6(a) and 6(b). The performance of CudaForge monotonically improves as the API and computation time increase. Notably, CudaForge already surpasses the Agentic baseline [2] while using no more than \$0.15 and 10 minutes per task, demonstrating its strong cost–performance tradeoff.
@@ 3.7 Generalization Capability of CudaForge
As shown in Figure 7, increasing $N$ from 1 to 10 leads to substantial performance gains, indicating that CudaForge can rapidly improve kernel efficiency through iterative refinement. Further increasing $N$ from 10 to 30 continues to improve performance, though with a slower growth rate, suggesting that the system gradually approaches its performance ceiling. These results demonstrate that CudaForge benefits from test-time scaling and has the potential to achieve even stronger performance given larger $N$ with additional inference cost.
@@ A.2 Prompt for Judge
In our prompt design for the Judge agent, we place the role specification and output schema in the system prompt. Inspired by [20, 21, 22], we design the prompts for correction and optimization, which are shown below.
@@ D.1 KernelBench
KernelBench has become a widely used benchmark in recent work on LLM-based code generation [16, 1, 2], as it provides a controlled and reproducible environment to measure both correctness (functional equivalence to PyTorch) and efficiency (execution speed relative to PyTorch). In our study, we adopt all Level 1–3 tasks, following prior work, to ensure fair comparison across baselines.

### v2

@@ Abstract
5555606065657070757580808585909095951001000.60.60.80.8111.21.21.41.41.61.61.81.8222.22.2OpenAI-$o3$-self-refineo3-correctiono3-optimizationCudaForgeCudaForgeCudaForge-Scaling UpKevin-32BAgentic BaselineCorrectness (%)Performance (× Speedup)w/o Judgew/o correction feedbackw/o optimization feedbackEvaluated on Level 1–2Evaluated on Level 1–3
[…]
@@ 1 Introduction
Motivation. CUDA has become the de facto standard for deep learning training because modern frameworks such as PyTorch and TensorFlow are deeply integrated with NVIDIA’s optimized GPU libraries [nvidia2025cudnn]. Efficient CUDA kernels are crucial for accelerating deep learning workloads [dao2022flashattention, dao2023flashattention2] .
However, developing high-efficiency CUDA kernels is known to be challenging with a steep learning curve, requiring deep expertise in GPU architectures and parallel programming [onegraph]. For example, it took more than 2 years from the debut of the Hopper GPU architecture to the release of FlashAttentionV3 [shah2024flashattention3fastaccurateattention], which is specially designed for Hopper GPUs.
This high development barrier has driven growing interest in finding automated ways of generating highly efficient and customized CUDA kernels. For example, some work [DBLP:conf/pldi/TilletKC19] [chen2019learningoptimizetensorprograms] employs auto-tuning and evolutionary search to automatically explore kernel implementation spaces and optimize low-level parameters for specific hardware. More recently, there has been a growing interest in leveraging large language models (LLMs) to perform such tasks. LLM is believed to hold great promise in generating efficient and high-quality kernels, due to its capability of code generation in other domains, such as Python and C++ [dong2025surveycodegenerationllmbased, jiang2024surveylargelanguagemodels].
Existing Works and Key Challenges. Generally, using LLMs for CUDA kernel generation is still in an early stage. In KernelBench [ouyang2025kernelbenchllmswriteefficient], the authors attempt to directly use state-of-the-art (SOTA) models, such as OpenAI-o1 and Claude-3.5-Sonnet, to generate kernels. However, it has been observed that these SOTA models still struggle to produce correct or performant kernels out of the box, revealing fundamental limitations of existing LLMs in this domain.
To address this gap, recent studies have explored two main paradigms. The first approach is based on reinforcement learning (RL) [schulman2017proximalpolicyoptimizationalgorithms, shao2024deepseekmathpushinglimitsmathematical]. CUDA-L1 [deepreinforce2025cudal1] and Kevin [kevin-multi-turn-rl] adopt RL to enhance LLMs’ ability to generate correct and performant CUDA code.
The second approach is based on AI agents. In particular, in an independent and contemporaneous work [lange2025robustagenticcudakernel]111published on arxiv Sept 16th, 2025, researchers have explored agentic frameworks at inference time. Agents project PyTorch methods into CUDA kernel designs, then the CUDA kernels are further refined by sampling new kernels and verification filtering. This design effectively improves correctness in CUDA kernel generation without the high cost of RL training.
@@
(C2) High training and inference cost. RL-based approaches such as [deepreinforce2025cudal1, kevin-multi-turn-rl] require substantial computational resources and long training cycles, making them unsuitable for low-resource or rapid-prototyping settings. In addition, multi-stage agentic pipeline developed by [lange2025robustagenticcudakernel] incurs high inference costs (about 6 H100 hours and \$5 API cost per kernel), which greatly limits its practical applicability of the approach.
(C3) Lack of hardware feedback. Human experts typically follow an iterative workflow to develop CUDA kernels through testing and refinement. They rely on hardware feedback like Nsight Compute (NCU) [^2] Nsight Compute (NCU) is NVIDIA’s official kernel-level profiler for CUDA programs. to identify bottlenecks and optimize kernels accordingly [10946796, cudaguide, w4a8gemm]. In contrast, RL-based approaches [deepreinforce2025cudal1, kevin-multi-turn-rl] train LLMs to directly generate or optimize kernels, but do not leverage hardware feedback at all. As a result, they rely on blind exploration during generation, lacking targeted guidance. This often leads to suboptimal kernel efficiency, limiting their practical applicability.
@@
Our Contributions. To address these challenges, we propose CudaForge, a simple, effective, and low-cost multi-agent workflow for CUDA kernel generation and optimization, as shown in Figure 2. Our workflow is inspired by the iterative workflow of human experts [10946796, cudaguide, w4a8gemm], which contains steps such as developing initial kernels, testing correctness, analyzing hardware feedback, and iterative improvement.
@@
Compared to single-LLM approaches that generate and evaluate code using the same LLM, our framework separates these roles into an independent Coder and Judge, enabling more specialized reasoning and more reliable iterative refinement. Unlike RL-based methods, CudaForge is training-free, avoiding the substantial cost of policy training. It is also hardware-aware, allowing it to tailor CUDA kernel optimizations to the underlying system, making the proposed framework easily generalizable across different GPUs. Finally, in contrast to existing multi-agent frameworks [lange2025robustagenticcudakernel], CudaForge is lightweight and cost-efficient, running in just 26.5 minutes on a single RTX 6000 GPU and about \$ 0.3 per kernel in API on average, while still achieving significantly better performance.
We evaluate CudaForge on 250 KernelBench tasks from Level 1 to Level 3. Though these tasks are challenging, CudaForge attains a 97.6% correctness rate and delivers an average speedup of 1.68× over PyTorch baselines, which significantly outperforms advanced RL model like Kevin-32B, advanced frontier model like OpenAI-o3 [openai2025o3systemcard] and Agentic Baseline [lange2025robustagenticcudakernel], shown in Figure 1. Further, we have conducted comprehensive ablation studies of the features of CudaForge, such as its effectiveness across multiple GPU architectures, its inference-time scalability by increasing the number of generation, and the effect of different base models. Overall, we observed that the proposed CudaForge achieves robust performance in all these settings.
@@ 2.1 CudaForge Framework
Inspired by the iterative workflow of human experts [10946796, cudaguide, w4a8gemm], we design CudaForge as an iterative multi-agent framework, illustrated in Figure 2. The framework involves two independent agents: a Coder and a Judge. The Coder generates candidate kernels based on the task description and feedback from the Judge, while the Judge evaluates each candidate using the kernel itself, hardware feedback, and runtime information.
@@ 2.2 Component Design
The Coder acts as the generative module in our framework, responsible for producing functionally correct and efficient CUDA kernel candidates given the task description, previous candidates, and feedback from the Judge. Directly generating a CUDA kernel from scratch is a highly challenging task for LLMs [ouyang2025kernelbenchllmswriteefficient]. Following the setting of KernelBench [ouyang2025kernelbenchllmswriteefficient], we provide the Coder with a one-shot demonstration in the prompt, which includes an example pair of a PyTorch reference and its corresponding CUDA kernel implementation. This example serves as a structural and stylistic guide, helping the Coder understand the expected syntax, memory access patterns, and API usage required for valid kernel generation.
@@
After the Coder generates a candidate kernel, our workflow proceeds to verify its correctness through a two-stage testing process, including compilation and execution. The compilation stage ensures that the generated kernel is syntactically valid and can be successfully compiled into executable CUDA code. The execution stage then evaluates the functional correctness of the compiled kernel by running it on predefined test cases. We compare the kernel’s outputs with those from the PyTorch reference implementation under the same inputs, and consider the kernel functionally correct only if the numerical difference between the two outputs is within a small tolerance (set to $1\text{e}{-4}$ in our experiments). A kernel is regarded as correct only when it passes both compilation and execution stages successfully, which is a commonly adopted criterion [kevin-multi-turn-rl, lange2025robustagenticcudakernel, ouyang2025kernelbenchllmswriteefficient].
@@ 3.1 Benchmark and Evaluation
We evaluate our method and baselines on KernelBench [ouyang2025kernelbenchllmswriteefficient], a popular benchmark designed to assess the ability of LLMs to generate CUDA kernels. KernelBench consists of multiple difficulty levels, and we adopt all tasks from Level 1 to Level 3, resulting in a total of 250 tasks. Specifically, Level 1 contains relatively simple 100 tasks involving basic operators (e.g., matrix multiplication), Level 2 includes medium-difficulty 100 tasks composed of multi-step operator combinations, and Level 3 contains 50 challenging tasks involving full neural network architectures (e.g., AlexNet). Each task is accompanied by a reference PyTorch implementation and predefined input/output specifications, which enables fully automated and reliable evaluation of both correctness and performance. Details of KernelBench are provided in Appendix D.1.
@@ 3.2 Settings & Baselines
  • Kevin-32B: a strong RL-based model for CUDA kernel optimization from [kevin-multi-turn-rl]. We directly take results from their official paper.
  • Agentic Baseline: the agentic workflow from [lange2025robustagenticcudakernel], a strong multi-agent baseline, which uses an LLM ensemble including both reasoning (o3 & o4-mini) and conventional LLMs (Claude Sonnet 3.7 & GPT-4.1).
@@
Table 1: Main results on KernelBench (Level 1-3, 250 tasks). Results of Agentic Baseline is on Level 1 and 2. All experiments here are run in RTX 6000. Methods evaluated on $\mathcal{D}^{*}$ are marked with . CudaForge-Scaling Up means scaling up maximum iteration rounds
@@ 3.4 Comparison with Kevin-32B on H200
In this subsection, we compare CudaForge with Kevin-32B [kevin-multi-turn-rl]. However, since Kevin-32B is not open-sourced—meaning that key details such as test-time prompts, evaluation setup, and benchmark specifications are not publicly available—a fully fair and reproducible comparison cannot be conducted. According to the original paper, Kevin-32B was trained and evaluated on H200 GPUs using a self-constructed benchmark of comparable difficulty to KernelBench Levels 1 and 2. To ensure the comparison is as fair as possible under these constraints, we re-evaluate CudaForge on the same H200 hardware across KernelBench Levels 1–3, aligning as much as possible our computational environment and evaluation protocol with those reported for Kevin-32B.
@@ 3.5 API and Computation Time Cost of CudaForge
Table 3: Comparison of API and computation time cost between CudaForge and the Agentic Baseline [lange2025robustagenticcudakernel].
@@
As shown in Table 3, CudaForge requires on average only 26.5 minutes of wall-clock time on a single RTX 6000 GPU and incurs merely \$0.3 of API cost per kernel. This is highly cost-efficient compared with the Agentic Baseline [lange2025robustagenticcudakernel], which reports approximately 6 GPU hours on H100 and \$5 per kernel in their Appendix E.
@@
We further analyze how the API and computation costs relate to CudaForge’s performance, as illustrated in Figure 6(a) and 6(b). The performance of CudaForge monotonically improves as the API and computation time increase. Notably, CudaForge already surpasses the Agentic baseline [lange2025robustagenticcudakernel] while using no more than \$0.15 and 10 minutes per task, demonstrating its strong cost–performance tradeoff.
@@ 3.7 Generalization Capability of CudaForge
[…]
@@ A.2 Prompt for Judge
In our prompt design for the Judge agent, we place the role specification and output schema in the system prompt. Inspired by [agashe2024agentsopenagentic, lei2025infantagentnextmultimodalgeneralistagent, chen2023agentversefacilitatingmultiagentcollaboration], we design the prompts for correction and optimization, which are shown below.
@@ D.1 KernelBench
KernelBench has become a widely used benchmark in recent work on LLM-based code generation [deepreinforce2025cudal1, kevin-multi-turn-rl, lange2025robustagenticcudakernel], as it provides a controlled and reproducible environment to measure both correctness (functional equivalence to PyTorch) and efficiency (execution speed relative to PyTorch). In our study, we adopt all Level 1–3 tasks, following prior work, to ensure fair comparison across baselines.